OpenStudy (anonymous):

The perimeter of a rectangle is 36 m. If its length is increased by 1 m and the width is increased by 2m, its area will increase by 30m^2. Find the area of the original rectangle.

OpenStudy (anonymous):

Opposite sides of a rectangle are equal. Let x = length and y = width. Knowing that the area of a rectangle is length times width and that the perimeter is 2x + 2y, P = 36 = 2x+ 2y xy + 30 = (x + 1)(y + 2) original area is 30 less than the new area, xy + 30 = xy + 2x + y + 2 30 = 2x + y + 2 28 = 2x + y 36 = 2x + 2y subtract the two equations to get -8 = -y y = 8 28 = 2x + 8 2x = 20 x = 10 Checking this, perimeter of a 10x8 rectangle is 20 + 16 = 36. Area would be 80. Adding one to the length and two to the width, the rectangle becomes a 11x10 rectangle. Area of that would be 110. Since 110 - 80 is 30, all of the constraints are satisfied; the answer is 80m²